OpenStudy (anonymous):

Which equation has a solution of d = 145?

OpenStudy (kewlgeek555):

Are you going to give out the answer options? Or does this assignment require you making one up?

OpenStudy (anonymous):

A 157 – d = 302 B 462 – d = 327 C 517 = d + 382 D 359 = d + 214

OpenStudy (nirmalnema):

d is the correct ans..

OpenStudy (nirmalnema):

put the value of d in each eq..and check out

OpenStudy (kewlgeek555):

kay. So what I would do is plug in the choices. A. 157 - 145 = 302? Nope. B. 462 - 145 = 327? Nope. C. 382 + 145 = 517? Nope. Then it must be D! But just to make sure: 214 + 145 = 359? Yep!